# Multi-Objective Reinforcement Learning for Adaptive Personalized Autonomous Driving

### v1

@@ Abstract
Human drivers exhibit individual preferences regarding driving style. Adapting autonomous vehicles to these preferences is essential for user trust and satisfaction. However, existing end-to-end driving approaches often rely on predefined driving styles or require continuous user feedback for adaptation, limiting their ability to support dynamic, context-dependent preferences. We propose a novel approach using multi-objective reinforcement learning (MORL) with preference-driven optimization for end-to-end autonomous driving that enables runtime adaptation to driving style preferences. Preferences are encoded as continuous weight vectors to modulate behavior along interpretable style objectives—including efficiency, comfort, speed, and aggressiveness—without requiring policy retraining. Our single-policy agent integrates vision-based perception in complex mixed-traffic scenarios and is evaluated in diverse urban environments using the CARLA simulator. Experimental results demonstrate that the agent dynamically adapts its driving behavior according to changing preferences while maintaining performance in terms of collision avoidance and route completion.  [^1] Code will be made publicly available upon paper acceptance.
@@ I Introduction
However, existing vision-based end-to-end AD approaches typically rely on static preference configurations using lexicographic MORL [7, 8, 9], or require extensive human expert input [10]. Others necessitate re-optimization whenever preferences change [11], or learn multiple policies to approximate preference trade-offs [12], all offering limited support for efficient and dynamic preference adaptation.
@@
To address this, we propose a single-policy, preference-driven multi-objective reinforcement learning approach for end-to-end AD. Preferences are encoded as a continuous weight vector and provided at runtime to modulate the agent’s behavior along different driving styles, including comfort, efficiency, speed, aggressiveness. Our setup allows post-training adaptation to diverse and changing user preferences, without retraining or maintaining multiple policies.
@@
  • We design a reward vector that separates core driving task from preference-based style objectives, enabling smooth transitions across the driving styles comfort, speed, aggressiveness and efficiency.
@@ II Related Work
Personalization in autonomous driving is increasingly recognized as essential for enhancing user experience, trust, satisfaction, and long-term acceptance by aligning vehicle behavior with individual needs [18, 6, 19]. Traditional approaches explored this alignment through control-based methods [20, 21]. Li et. al. [22] embed personalized constraints into a Model Predictive Control (MPC) framework extended with proportional–integral–derivative (PID) feedback to reflect individual driving styles via explicit control parameter tuning. However, such methods depend on handcrafted control rules and are often limited to specific scenarios, lacking the scalability of learning-based approaches.
@@
Multi-objective reinforcement learning extends standard RL to vector-based reward functions with policies that learn to balance competing objectives, making it particularly well-suited for preference-driven end-to-end AD. Li and Czarnecki [8] apply thresholded lexicographic Q-learning to urban driving, optimizing fixed objectives such as safety and comfort, each managed by a separate agent, but without vision-based inputs as addressed in our approach. Hu et al. [9] propose a MORL framework that balances speed and comfort in highway-like scenarios using a Unity-based simulator. However, their setup is limited by simplified dynamics utilizing discrete high-level actions, as well as by only two objectives and fixed preferences. In contrast, our approach supports dynamic preference integration in complex urban environments with a broader range of personalized driving styles. Wang et al. [26] demonstrates the potential of MORL for end-to-end AD by learning Pareto-optimal policies over objectives like speed and fuel efficiency using image-based input. However, the approach optimizes only a single preference configuration per training run, requiring retraining for new preferences and thus limiting runtime adaptability.
Recent single-policy MORL frameworks show promise for runtime adaptability of preference [27, 28], but have not yet been applied to vision-based end-to-end autonomous driving. We extends that line by integrating an adaptive, preference driven framework into high-dimensional, vision-based urban driving scenarios, as illustrated in Figure 2.
@@ III-A Preference-Based Driving Task
We consider an RL-driven autonomous vehicle navigating through dense urban environments with dynamic traffic under varying weather and lighting conditions, trained and evaluated with the CARLA simulator [29]. The agent’s objective is to navigate a predefined route represented by waypoints, while avoiding collisions and adapting its behavior to given human driving preferences. These preferences—aggressiveness, comfort, speed and efficiency—may vary with time or task and are encoded as a four-dimensional vector to modulate the agent’s behavior. The agent receives semantic segmentation (SEG) from a front-facing camera as input and has access to the vehicles odometry, the current speed limit, and the next waypoint. While it must comply with the current speed limit, advanced traffic rules are not considered. The vehicle is controlled via continuous steering and throttle/brake commands as direct output from the agent.
@@ III-B Multi-Objective Reinforcement Learning
Multi-objective reinforcement learning (MORL) extends traditional reinforcement learning by optimizing multiple, often conflicting objectives [30]. In MORL, the agent is tasked with learning policies that reflect trade-offs across these objectives by utilizing a vector-valued reward function. Formally, the learning problem is defined as a Multi-Objective Markov Decision Process (MO-MDP), specified by the tuple $(\mathcal{S},\mathcal{A},P,\mathcal{R},\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ the transition dynamics, and $\gamma\in[0,1]$ the discount factor. A key component is the vector-valued reward function $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{n}$, which returns $n$ objective-specific rewards at each time step. A single policy $\pi^{\boldsymbol{\lambda}}$ can be conditioned on a user-defined preference vector $\boldsymbol{\lambda}\in\mathbb{R}^{n}$ to express trade-offs between multiple objectives. The preferences are incorporated into the learning process via element-wise multiplication with the Q-values $\boldsymbol{\lambda}\odot\mathbf{Q}_{j}$.
@@ III-C State and Action Space
Vehicle odometry $\boldsymbol{o}$ includes velocity and acceleration (both magnitudes and directional components), along with the current positions of the throttle, brake, and steering controls. Temporal context is added by including the previous three control inputs $\boldsymbol{a}_{\text{hist}}$ in the state, as vehicle behavior exhibits temporal dependencies, enabling the agent to learn vehicle dynamics more accurately.
@@ III-F Reward Vector
The reward vector $\boldsymbol{r_{t}}$ captures both the core driving task and user-dependent driving style objectives. The core driving reward occupies the first entry in vector and captures essential driving and navigation competence. The remaining four components correspond to distinct driving behavioral preferences: aggressiveness, comfort, speed and efficiency.
@@
Figure 3: a) Bird’s-eye view of the agent’s trajectories while navigating a T-intersection. Each trajectory reflects a distinct driving style, achieved by setting one preference weight (for speed, efficiency, aggressiveness, or comfort) to one, with all others set to zero. Subfigures b) - e) display additional metrics including steering, throttle, velocity, and lateral acceleration. Values are smoothed using exponential smoothing with a factor of $\beta=0.6$. Start and end points are denoted by ”X” and $\blacktriangle$”, respectively.
@@ III-F1 Core Driving Reward
The static core reward remains unaffected by user preferences and captures fundamental driving capabilities. It encourages route progression, while penalizing unsafe behaviors such as collisions, off-road driving, lane invasions, and speed limit violations, braking in hazardous situations, and appropriate velocity. The core reward is therefore composed of five components:
@@
The detailed formulation of each component is provided in the Appendix -A. Combined, these elements ensure stable and safe driving behavior independent of user preference modulation.
@@ III-F2 Preference-Based Objectives
The preference-based reward vector $\boldsymbol{r}_{\text{pref}}$ encodes the driving style objectives, with each dimension corresponding to a specific preference.
@@
$\cdot|\Delta v|-\beta_{\text{long}}\cdot|\Delta a_{\text{long}}|-\beta_{\text{jerk}}\cdot||\text{jerk}||+\beta_{\text{b}}.$
Speed ($r_{\text{speed}}$): Rewards matching a high target velocity $v_{\text{target}}$ based on the speed limit. Deviations are penalized as:
$r_{\text{speed}}=-\delta_{speed}\cdot\left(\frac{|v_{\text{target}}-v_{\text{current}}|}{v_{\text{target}}}\right)+\delta_{\text{speed}}.$
Efficiency ($r_{\text{eff}}$): Encourages fuel-efficient driving by minimizing throttle and acceleration, and promoting moderate speeds. Approximated as:
@@ IV-A Qualitative Preference Analysis
Figure 3 shows a qualitative analysis of the agent’s driving behavior under different preference objectives, presenting both its spatial trajectories and corresponding graphs illustrating its behavior across driving metrics.
@@
The results show that the policy adapts to different preference configurations, displaying distinct and interpretable driving styles.
@@ IV-B1 Preference Reflection
Each subplot shows the mean metric value across episodes, with solid lines denoting the average trend and shaded regions indicating one standard deviation. The x-axis represents the fixed sampled weight of the respective objective, shared across all subplots.
@@
In support of RQ1, the results confirm that varying the preference weightings induces significant shifts in driving behavior, demonstrating the policy’s responsiveness to various preferences.
@@ V Conclusion
In summary, we present a multi-objective reinforcement learning approach capable of integrating dynamic user preferences into end-to-end vision-based autonomous driving in the CARLA simulator. Our method employs a single policy conditioned on a continuous, user-defined preference vector, enabling runtime adaptation of driving behavior without retraining. Leveraging vision-based perception, the agent learns to modulate its driving style in response to varying preference weightings. Experimental evaluations in complex urban traffic scenarios demonstrate that the agent maintains robust driving performance while reflecting user preferences across diverse metrics. To improve perception quality, we evaluate multiple visual input modalities and feature extraction architectures, with segmentation input and a compact ResNet-based encoder achieving the best overall performance.

### v2

# Multi-Objective Reinforcement Learning for Adaptable Personalized Autonomous Driving

@@ Abstract
Human drivers exhibit individual preferences regarding driving style. Adapting autonomous vehicles to these preferences is essential for user trust and satisfaction. However, existing end-to-end driving approaches often rely on predefined driving styles or require continuous user feedback for adaptation, limiting their ability to support dynamic, context-dependent preferences. We propose a novel approach using multi-objective reinforcement learning (MORL) with preference-driven optimization for end-to-end autonomous driving that enables runtime adaptation to driving style preferences. Preferences are encoded as continuous weight vectors to modulate behavior along interpretable style objectives—including efficiency, comfort, speed, and aggressiveness—without requiring policy retraining. Our single-policy agent integrates vision-based perception in complex mixed-traffic scenarios and is evaluated in diverse urban environments using the CARLA simulator. Experimental results demonstrate that the agent dynamically adapts its driving behavior according to changing preferences while maintaining performance in terms of collision avoidance and route completion. Our code is available online on https://github.com/HumanoidsBonn/morl_preference_driving.
@@ I Introduction
However, existing vision-based end-to-end AD approaches typically rely on static preference configurations using lexicographic multi-objective reinforcement learning (MORL) [7, 8, 9], or require extensive human expert input [10]. Others necessitate re-optimization whenever preferences change [11], or learn multiple policies to approximate preference trade-offs [12], all offering limited support for efficient and dynamic preference adaptation.
@@
To address this, we propose a single-policy, preference-driven multi-objective reinforcement learning approach for end-to-end AD. Preferences are encoded as a continuous weight vector and provided at runtime to modulate the agent’s behavior along different driving styles, namely comfort, efficiency, speed, and aggressiveness. Our setup allows post-training adaptation to diverse and changing user preferences, without retraining or maintaining multiple policies.
@@
  • We design a reward vector that separates the core driving task from preference-based style objectives, enabling smooth transitions across the driving styles comfort, speed, aggressiveness and efficiency.
@@ II Related Work
Personalization in autonomous driving is increasingly recognized as essential for enhancing user experience, trust, satisfaction, and long-term acceptance by aligning vehicle behavior with individual needs [6, 18, 19]. Traditional approaches explored this alignment through control-based methods [20, 21]. Li et al. [22] embed personalized constraints into a Model Predictive Control (MPC) framework extended with proportional–integral–derivative (PID) feedback to reflect individual driving styles via explicit control parameter tuning. However, such methods depend on handcrafted control rules and are often limited to specific scenarios, lacking the scalability of learning-based approaches.
@@
Multi-objective reinforcement learning extends standard RL to vector-based reward functions with policies that learn to balance competing objectives, making it particularly well-suited for preference-driven end-to-end AD. Li and Czarnecki [8] apply thresholded lexicographic Q-learning to urban driving, optimizing fixed objectives such as safety and comfort, each managed by a separate agent, but without vision-based inputs as addressed in our approach. Hu et al. [9] propose a MORL framework that balances speed and comfort in highway-like scenarios using a Unity-based simulator. However, their setup is limited by simplified dynamics utilizing discrete high-level actions, as well as by only two objectives and fixed preferences. In contrast, our approach supports dynamic preference integration in complex urban environments with a broader range of personalized driving styles. Wang et al. [26] demonstrate the potential of MORL for end-to-end AD by learning Pareto-optimal policies over objectives like speed and fuel efficiency using image-based input. However, the approach optimizes only a single preference configuration per training run, requiring retraining for new preferences and thus limiting runtime adaptability.
Recent single-policy MORL frameworks show promise for runtime adaptability of preferences [27, 28], but have not yet been applied to vision-based end-to-end autonomous driving. We extend that line by integrating an adaptable, preference driven framework into high-dimensional, vision-based urban driving scenarios, as illustrated in Figure 2.
@@ III-A Preference-Based Driving Task
We consider an RL-driven autonomous vehicle navigating through dense urban environments with dynamic traffic under varying weather and lighting conditions, trained and evaluated with the CARLA simulator [29]. The agent’s objective is to navigate a predefined route represented by waypoints, while avoiding collisions and adapting its behavior to given human driving preferences. These preferences—aggressiveness, comfort, speed and efficiency—may vary with time or task and are encoded as a four-dimensional vector to modulate the agent’s behavior. The agent receives semantic segmentation (SEG) from a front-facing camera as input and has access to the vehicle’s odometry, the current speed limit, and the next waypoint. While it must comply with the current speed limit, advanced traffic rules are not considered. The vehicle is controlled via continuous steering and throttle/brake commands as direct output from the agent.
@@ III-B Multi-Objective Reinforcement Learning
MORL extends traditional reinforcement learning by optimizing multiple, often conflicting objectives [30]. The agent is tasked with learning policies that reflect trade-offs across these objectives by utilizing a vector-valued reward function. Formally, the learning problem is defined as a Multi-Objective Markov Decision Process (MO-MDP), specified by the tuple $(\mathcal{S},\mathcal{A},P,\mathcal{R},\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ the transition dynamics, and $\gamma\in[0,1]$ the discount factor. A key component is the vector-valued reward function $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{n}$, which returns $n$ objective-specific rewards at each time step. A single policy $\pi^{\boldsymbol{\lambda}}$ can be conditioned on a user-defined preference vector $\boldsymbol{\lambda}\in\mathbb{R}^{n}$ to express trade-offs between multiple objectives. The preferences are incorporated into the learning process via element-wise multiplication with the Q-values $\boldsymbol{\lambda}\odot\mathbf{Q}_{j}$.
@@ III-C State and Action Space
Vehicle odometry $\boldsymbol{o}$ includes velocity and acceleration (both magnitudes and directional components), along with the current positions of the throttle, brake, and steering controls. By incorporating the previous three control inputs $\boldsymbol{a}_{\text{hist}}$ into the state, temporal dependencies in vehicle behavior are captured, assisting the agent in learning smooth vehicle dynamics [32, 33, 34].
@@ III-F Reward Vector
The reward vector $\boldsymbol{r_{t}}$ captures both the core driving task and user-dependent driving style objectives. The core driving reward occupies the first entry in the vector and captures essential driving and navigation competence. The remaining four components correspond to distinct driving behavioral preferences: aggressiveness, comfort, speed and efficiency.
@@
Figure 3: a) Bird’s-eye view of the agent’s trajectories while navigating a T-intersection. Each trajectory reflects a distinct driving style, achieved by setting one preference weight (for speed, efficiency, aggressiveness, or comfort) to one, with all others set to zero. Subfigures b) - e) display additional metrics including steering, throttle, velocity, and lateral acceleration. Values are smoothed using exponential smoothing with a factor of $\beta=0.6$.
@@ III-F1 Core Driving Reward
The static core reward remains unaffected by user preferences and captures fundamental driving skills. It encourages route progression, braking in hazardous situations, and maintaining an appropriate velocity, while penalizing unsafe behaviors, including collisions, off-road driving, lane invasions, and speed limit violations [36, 37]. The core reward is composed of five components:
@@
The detailed formulation of each component is provided in the Appendix -A. Combined, these elements ensure stable and safe driving behavior regardless of user preference modulation.
@@ III-F2 Preference-Based Objectives
The preference-based reward vector $\boldsymbol{r}_{\text{pref}}$ encodes driving style objectives, with each dimension corresponding to a specific preference, guided by concepts from [26, 36, 38].
@@
$\cdot|\Delta v|-\beta_{\text{long}}\cdot|\Delta a_{\text{long}}|-\beta_{\text{jerk}}\cdot\|\text{jerk}\|+\beta_{\text{b}}.$
Speed ($r_{\text{speed}}$): Rewards matching the target velocity $v_{\text{target}}$ based on the speed limit, with deviations from the target velocity penalized as:
$r_{\text{speed}}=-\delta_{\text{speed}}\cdot\left(\frac{|v_{\text{target}}-v_{\text{current}}|}{v_{\text{target}}}\right)+\delta_{\text{speed}}.$
Efficiency ($r_{\text{eff}}$): Encourages fuel-efficient driving by minimizing throttle as well as acceleration, and promoting moderate speeds. It is approximated by:
@@ IV-A Qualitative Preference Analysis
Figure 3 shows a qualitative analysis of the agent’s driving behavior under different preference vectors, presenting both top-down view trajectories and corresponding driving metric graphs.
@@
The results show that the policy adapts to different preference configurations, displaying distinct and interpretable driving styles. A supplemental video illustrating the agent’s qualitative behavior is available online. [^1] Supplemental video: https://youtu.be/2brpyC_edHw
@@ IV-B1 Preference Reflection
Each subplot shows a specific driving metric relevant to one or more driving objectives, with mean (solid line) and standard deviation (shaded area) calculated across evaluation episodes. The selected preferences in each subplot correspond to their conceptually matching driving metric. The x-axis represents the fixed sampled weight of the respective objective, shared across all subplots.
@@
In support of RQ1, the results confirm that varying the preference weights induces significant shifts in driving behavior, demonstrating the policy’s responsiveness to various preferences.
@@ V Conclusion
In summary, we present a multi-objective reinforcement learning approach capable of integrating dynamic user preferences into end-to-end vision-based autonomous driving in the CARLA simulator. Our method employs a single policy conditioned on a continuous, user-defined preference vector, enabling runtime adaptation of driving behavior without retraining. Leveraging vision-based perception, the agent learns to modulate its driving style in response to varying preference weightings. Experimental evaluations in complex urban traffic scenarios demonstrate that the agent maintains robust driving performance while reflecting user preferences across diverse metrics. To improve perception quality, we evaluate multiple visual input modalities and feature extraction architectures, with segmentation input and a compact ResNet-based encoder achieving the best overall performance. Future work could extend this framework by modeling more sophisticated traffic rules and validating it on real-world driving platforms.